# Ruelle-Pollicott resonances of diffusive U(1)-invariant qubit circuits

### v1

@@ III.1 Small quasi-momenta and magnetization transport
Figure 5: Leading RP resonance and correlation function decay in the $k=\pi$ spectrum for magnetization-conserving circuits. (a) The whole spectrum of the truncated propagator in $k=0,\pi$ sectors for case 3 and $r=7$. (b) Convergence of the leading RP resonances in the $k=\pi$ sector with $r$, with an extrapolation to the $r\to\infty$ limit. (c) Infinite temperature autocorrelation functions of extensive observables $A$ defined in Eq. (5) with $k=\pi$. The local densities are either $\sigma^{x}$ (i.e., $x$ magnetization) or chosen randomly according to the Gaussian unitary ensemble Mehta (2004). The correlation functions are calculated in a finite circuit with $N=30$ sites. Dashed lines show the RP predictions $\propto\left\lvert\lambda_{1}\right\rvert^{t}$.
Both methods of determining $D$ are shown in Fig. 4. In the transferred magnetization from the domain-wall (bottom left (ii) insets), we again see $z$ matching the expected $z=2$. The diffusion constants determined from different approaches also match up to accessible precision. We have thus shown the leading RP resonance for small $k$ truly does contain information about transport. A natural next question, which we leave for future work, is how competitive is the extraction of $D$ through RP resonances compared to the current state-of-the-art.
III.2 Quasi-momenta far from zero and correlation function decay
@@
The leading RP resonances far from $k=0$ do not follow ${\rm e}^{-Dk^{2}}$ anymore and are not connected to diffusive transport, see Fig. 3b. We thus expect them to merely be indicators of exponential decay, like in the case without symmetries in Fig. 2. We demonstrate this numerically for $k=\pi$ in Fig. 5. Convergence with $r$ is slower than at small $k$ and at the largest available $r=13$ the leading eigenvalue has not yet fully converged, however, it does seem that the extrapolated values  [^3] The extrapolation is done through an exponential fit. We have no good theoretical justification for the exponential fit, it was simply the best match for the data. are strictly smaller than $1$. Therefore, generic correlation functions are expected to decay exponentially with a finite rate. In Fig. 5c we show correlation functions of $x$ magnetization and of a random observable. While exponential decay at available system size $N=30$ is not super nice, the decay is compatible with the calculated extrapolated $\lambda_{1}$, e.g., it is faster for the circuit with smaller $\left\lvert\lambda_{1}\right\rvert$. Also, different observables decay with the same rate.
@@
The fact that the exponential decay in the present case is not as evident from numerics is compatible with slow convergence of the RP resonance with $r$. Namely, the leading RP resonance at truncation $r$ can be intuitively understood to govern the decay of correlations when the operator has spread to $r$ sites. Slower convergence of RP resonances with $r$ thus implies “nice” exponential decay only at later times and consequently smaller values of $C_{AA}$. Therefore, it is expected that exponential decay will be harder to see numerically at finite $N$ precisely when RP resonances will also be hard to extract numerically. Additionally, correlation functions in Fig. 5c also visibly oscillate, making numerical checks less clear. This, however, is merely a consequence of $\lambda_{1}$ having an imaginary component seen for the example of case 3 in Fig. 5a.
IV Beyond the leading Ruelle-Pollicott resonance
@@
We have shown that the leading RP resonance $\lambda_{1}(k)$ for small $k$ is connected to transport of the conserved quantity. How about subleading eigenvalues; they could contain information about other correlation functions. One might expect correlation functions of observables orthogonal to all transport-related quantities decay exponentially, like in a system without any conserved quantities, which would imply that there must exist subleading RP resonances.
@@
Figure 6: 20 leading eigenvalues of the truncated propagator at different $k$. Calculated for case 3 and support $r=10$.
@@
The first few largest eigenvalues of the truncated propagator can be calculated numerically and are shown in Fig. 6. We observe many subleading eigenvalues in a narrow interval and one might think that they are isolated RP resonances. It turns out this is not true; these eigenvalues have not yet converged with $r$ and, moreover, many of them look to be converging to $\lambda=1$. We shall propose in Sec. IV.1 that such eigenvalues close to $1$ come from the powers of the conserved magnetization $M$. Furthermore, we shall argue in Sec. IV.2 that power-law hydrodynamic tails in transport-related correlation functions imply an RP continuum in the spectrum of the truncated propagator as sketched in Fig. 1b. Both of these results make it hard to numerically determine whether subleading discrete RP resonances exist.
IV.1 Powers of the conserved quantity
@@
Without truncation $M^{2}$ is conserved. This can be also seen explicitly, ${\mathcal{U}}(M^{2})=\sum_{i,k}{\mathcal{U}}(\sigma_{i}^{z}){\mathcal{U}}(\sigma_{k}^{z})$, which, after invoking definition of the spin current $j_{k}$ [see Eq. (21)], ${\mathcal{U}}(\sigma_{k}^{z})=\sigma_{k}^{z}+j_{k-1}-j_{k}$, can be rewritten as $\mathcal{U}(M^{2})=\sum_{i,k}(\sigma_{i}^{z}+j_{i-1}-j_{i})(\sigma_{k}^{z}+j_{k-1}-j_{k})=M^{2}$. Conservation comes about because the many-qubit terms $j_{i}j_{k}$ (and likewise $\sigma_{i}^{z}j_{k}$) sum to 00 due to alternating signs stemming from the continuity equation. Importantly, the cancellations occur only between terms with distance $\left\lvert k-i\right\rvert$ differing by $1$, i.e., the term $j_{i}j_{k}$ can only be generated by the propagation of $\sigma^{z}_{i}\sigma^{z}_{k}$, $\sigma^{z}_{i+1}\sigma^{z}_{k}$, $\sigma^{z}_{i}\sigma^{z}_{k+1}$ and $\sigma^{z}_{i+1}\sigma^{z}_{k+1}$ (and similarly for $\sigma_{i}^{z}j_{k}$).
@@
In the basis truncated to support $r$, some of the terms needed for cancellation are not included, resulting in an almost-conservation. Let us denote the truncated $M^{2}$ by $Q^{(r)}$, explicitly $Q^{(r)}=\sum_{i,k,|k-i|<r}\sigma_{i}^{z}\sigma_{k}^{z}$. In particular, the $j_{i}j_{k}$ terms stemming from the propagation of $\sigma_{i}^{z}\sigma_{k}^{z}$ with $\left\lvert k-i\right\rvert=r-1$, that would cancel out with the same (but negative) terms stemming from $\sigma^{z}_{i}\sigma^{z}_{k+1}$, do not cancel out, since $\sigma^{z}_{i}\sigma^{z}_{k+1}$ is not included in $Q^{(r)}$ (and similarly for $\sigma_{i}^{z}j_{k}$ terms). Importantly, for $\left\lvert k-i\right\rvert<r-1$ the cancellation still occurs and all such terms are conserved. The upshot of this is that $\mathcal{U}(Q^{(r)})$ does not conserve only the $\sigma_{i}^{z}\sigma_{k}^{z}$ terms with distance $\left\lvert k-i\right\rvert=r-1$, the number of which does not scale with $r$.
The number of terms in $Q^{(r)}$ is equal to the number of indices with $|k-i|<r$, i.e., it scales as $\sim r$ ($Nr$ in a system of length $N$). The number of the conserved terms in $\mathcal{U}(Q^{(r)})$ scales as $\sim r-c$, where $c$ is some constant not scaling with $r$. In the limit $r\to\infty$, the two are almost the same, and one can estimate that $Q^{(r)}$ is an almost eigenvector with eigenvalue
@@
$\lvert\lambda\rvert\approx\frac{r-c}{r}=1-\frac{c}{r}.$ (20)
@@
IV.2 Hydrodynamic tails
In systems with conserved quantities certain correlation functions related to transport exhibit power-law decays, i.e., the so-called hydrodynamic tails Forster (1990); Mukerjee et al. (2006), due to slow power-law spreading of inhomogeneities of the conserved density (distance and time scaling as $x^{z}\sim t$). A typical example is the correlation function of local magnetization (15), i.e., the Green’s function of the diffusion equation, decaying as $\sim 1/t^{1/2}$ at $x=0$. Similar power-law tails occur also in higher-order corrections to diffusion equation Forster (1990); Mukerjee et al. (2006), more recently studied in the context of effective field theories Doyon (2022); Matthies et al. (2024); Crossley et al. (2017); Michailidis et al. (2024).
Power-law decay $C(t)\sim t^{-\alpha}$, for some $\alpha>0$, being slower than exponential, cannot be governed by a discrete RP resonance. If we nonetheless assume it expands over a spectrum of eigenvalues as $C(t)=\sum_{j}c_{j}\lambda_{j}^{t}$, where $c_{j}$ are expansions coefficients [cf. Eq. (11)], it must expand over a continuum of eigenvalues (an “RP continuum”) with closing gap to $\lambda=1$  [^5] The argument can be illustrated more explicitly by writing the decomposition as an integral $C(t)=\sum_{j}c_{j}{\rm e}^{-\nu_{j}t}=\int_{0}^{\infty}c(\nu){\rm e}^{-\nu t}\,\mathrm{d}\nu$, where $\lambda={\rm e}^{-\nu}$, $c(\nu)$ is the continuous version of the expansion coefficients and we assumed $\lambda$ to be real for simplicity. In other words, $C(t)$ is the Laplace transform of $c(\nu)$. By employing the inverse Laplace transform, we see that $C(t)=t^{-\alpha}$ is governed by $c(\nu)=\frac{1}{\Gamma(\alpha)}\nu^{\alpha-1}\theta(\nu)$, where $\Gamma$ is the Gamma function and $\theta$ is the Heaviside step function. We observe that $c(\nu)\neq 0$ for all $\nu>0$ an RP continuum over all $\lambda\in[0,1]$.. This is precisely what we observed for the correlation function of magnetization in momentum space (16), with the gap closing at $k=0$. Additionally, the quasi-momentum dependence of the RP continuum was exactly determined in that case, $\left\lvert\lambda(k)\right\rvert={\rm e}^{-Dk^{2}}$. The expansion over eigenvalues for all $k$ is expected only for correlation functions of local observables. For extensive observables, the whole continuum must lie in one $k$ sector. A potential RP continuum governing a power-law decay of a $k=0$ extensive observable is depicted in the diagram in Fig. 8.
@@ IV.2 Hydrodynamic tails
Based on the above arguments we conjecture that there is an RP continuum of eigenvalues for all $|\lambda|$ below the leading RP resonance governing the transport of magnetization. In other words, all $|\lambda|\leq{\rm e}^{-Dk^{2}}$ are part of an RP continuum.
IV.2.1 Extensive spin current
@@
where $j_{i}$ is the current operator between sites $i$ and $i+1$, and we introduced the magnetization acting on sites $i$ to $l$ as $M_{[i,l]}\coloneq\sum_{j=i}^{l}\sigma^{z}_{j}$. The closed-form expression for $j_{i}$ is complicated to derive analytically for our $3-$layer brickwall circuit, but can be obtained numerically for a given gate by evaluating the LHS of Eq. (21) in a big enough finite circuit. Current densities $j_{i}$ are different for different values of $i$ modulo $3$, and have support on at most $9$ sites. For a 2-layer brickwall circuit (e.g., Fig. 2a) the expressions are more manageable (2 or 4 site operators) and are written out for any magnetization preserving gate in Ref. Žnidarič (2025a).
The asymptotics of the local spin current correlation function can be deduced from the local magnetization correlation function and the continuity equation. The diffusive local magnetization correlation function (15) (i.e., the diffusion equation Green’s function) implies
@@
because it appears in the Green-Kubo Kubo et al. (1991) equation for the diffusion constant  [^7] For discrete-time dynamics, the Green-Kubo formula reads $D=\lim_{t_{\mathrm{max}}\to\infty}\lim_{N\to\infty}\frac{1}{N}\left(\frac{1}{2}\left\langle J(0)J(0)\right\rangle+\sum_{t=1}^{t_{\mathrm{max}}}\left\langle J(t)J(0)\right\rangle\right)$. In the considered circuits, it gives $D\approx 0.31$ in case 1 and $D\approx 0.44$ in case 2, which matches our predictions with other methods up to accessible precision (see Fig. 4).. Using Eq. (22) and the fact that the equilibrium correlation function is translationally invariant in space, we get for the extensive current in the thermodynamic limit
@@
The prediction coming from the diffusive (i.e., first order) local magnetization correlation function, thus gives exactly 00, which can be traced back to the fact that $\left\langle j(x,t)j(0,0)\right\rangle$ is a total derivative in $x$ [see Eq. (22)]. This is very interesting as it means that the leading behavior of the correlation function of the extensive current is given by subleading corrections to the local magnetization correlation function. Autocorrelation function of $J$ is therefore a sensitive probe of subleading corrections to diffusion that would be perhaps hard to see in other correlation functions that are nonzero already in the leading order.
Figure 9: Infinite temperature autocorrelation of extensive spin current. The correlation functions are calculated in a case 3 finite circuit with different number of sites $N$. Dashed lines show power-law fits.
Numerical calculation of the extensive current correlation function for one case is shown in Fig. 9. It decays as a power law for any finite system size $N$, however, the power increases with $N$. It is not clear what will happen in the thermodynamic limit; its power-law dependence may converge to some finite power or the decay may become exponential.
In addition to the inconclusive numerics, it is also not clear what is theoretically expected. Power-law decay $\sim t^{-3/2}$ is expected in systems with an additional conserved quantity, e.g., in Hamiltonian systems Mukerjee et al. (2006), and can be explained by cross transport of the conserved quantities. In systems with only one U(1) conserved quantity a power-law decay with a higher power was predicted Michailidis et al. (2024), although exponential decay is also sometimes observed, e.g., in the kicked Ising model Yi-Thomas et al. (2024).

### v2

@@ 3.2 Quasi-momenta far from zero and correlation function decay
Figure 5: Leading RP resonance and correlation function decay in the $k=\pi$ sector for magnetization-conserving circuits. (a) The whole spectrum of the truncated propagator in $k=0,\pi$ sectors for the circuit with gate 3 and $r=7$. (b) Convergence of the leading RP resonances in the $k=\pi$ sector with $r$, with dashed lines denoting an extrapolation to the $r\to\infty$ limit. (c) Infinite temperature autocorrelation functions of extensive observables AA defined in Eq. (5) with $k=\pi$. The local densities are either $\sigma^{x}$ (i.e., $x$ magnetization) or chosen randomly according to the Gaussian unitary ensemble (GUE) [27] with support $r=2$. The correlation functions are calculated in a finite circuit with $N=30$ sites. Dashed lines show the RP predictions $\propto\left\lvert\lambda_{1}\right\rvert^{t}$.
@@
The leading RP resonances far from $k=0$ do not follow ${\rm e}^{-Dk^{2}}$ anymore and are not connected to diffusive transport, see Fig. 3b. We thus expect them to merely be indicators of exponential decay, like in the case without symmetries in Fig. 2. We demonstrate this numerically for $k=\pi$ in Fig. 5. Convergence with $r$ is slower than at small $k$ and at the largest available $r=13$ the leading eigenvalue has not yet fully converged, however, it does seem that the extrapolated values [^3] The extrapolation is done through an exponential fit. We have no good theoretical justification for the exponential fit, it was simply the best match for the data. are strictly smaller than $1$. Therefore, generic correlation functions are expected to decay exponentially with a finite rate. In Fig. 5c we show correlation functions of $\sigma^{x}$ magnetization and of a random 2-site observable. While exponential decay at available system size $N=30$ is not super nice, the decay is compatible with the extrapolated $\lambda_{1}$, e.g., it is faster for the circuit with smaller $\left\lvert\lambda_{1}\right\rvert$. Also, different observables decay with the same rate.
@@
The fact that the exponential decay in the shown circuits is not as evident from numerics is compatible with the slow convergence of the RP resonance with $r$. Namely, the leading RP resonance at truncation $r$ can be intuitively understood to govern the decay of correlations when the operator has spread to $r$ sites. Slower convergence of RP resonances with $r$ thus implies “nice” exponential decay only at later times and consequently smaller values of $C_{AA}$. Therefore, it is expected that exponential decay will be harder to see numerically at finite $N$ precisely when RP resonances will also be hard to extract numerically. Additionally, correlation functions in Fig. 5c visibly oscillate, making numerical checks less clear. This, however, is merely a consequence of $\lambda_{1}$ having an imaginary component, see, e.g., gate 3 in Fig. 5a.
4 Beyond the leading Ruelle-Pollicott resonance
@@
We have shown that the leading RP resonance $\lambda_{1}(k)$ for small $k$ is connected to transport of the conserved quantity. How about subleading eigenvalues; they could contain information about other correlation functions. One might expect that correlation functions of observables orthogonal to all transport-related quantities decay exponentially, like in a system without any conserved quantities, which would imply that there must exist subleading RP resonances.
@@
Figure 6: 20 leading eigenvalues of the truncated propagator at different $k$. Calculated for the circuit with gate 3 and support $r=11$. The subleading eigenvalues are not yet converged at accessible $r$, see the discussion in the main text and Fig. 7 for details.
@@
The first few largest eigenvalues of the truncated propagator can be calculated numerically and are shown in Fig. 6. We observe many subleading eigenvalues in a narrow interval and one might think that they are isolated RP resonances. It turns out this is not the case; these eigenvalues have not yet converged with $r$ and, moreover, many of them look to be converging to the leading eigenvalue $\lambda_{1}$ (see Fig. 7 for the convergence to $\left\lvert\lambda_{1}\right\rvert=1$ in the $k=0$ sector). We shall propose in Sec. 4.1 that in the $k=0$ sector, some of such eigenvalues close to $1$ come from the powers of the conserved magnetization $M$. Furthermore, we shall argue in Sec. 4.2 that power-law hydrodynamic tails in transport-related correlation functions imply an RP continuum in the spectrum of the truncated propagator as sketched in Fig. 1b. Both of these results make it hard to numerically determine whether subleading discrete RP resonances exist.
4.1 Powers of the conserved quantity
@@
Without truncation, $M^{2}$ is conserved. This can be also seen explicitly, ${\mathcal{U}}(M^{2})=$ $\sum_{i,k}{\mathcal{U}}(\sigma_{i}^{z}){\mathcal{U}}(\sigma_{k}^{z})$, which, after invoking definition of the spin current $j_{k}$ [see Eq. (21)], ${\mathcal{U}}(\sigma_{k}^{z})=\sigma_{k}^{z}+j_{k-1}-j_{k}$, can be rewritten as $\mathcal{U}(M^{2})=\sum_{i,k}(\sigma_{i}^{z}+j_{i-1}-j_{i})(\sigma_{k}^{z}+j_{k-1}-j_{k})=$ $M^{2}$. Conservation comes about because the many-qubit terms $j_{i}j_{k}$ (and likewise $\sigma_{i}^{z}j_{k}$) sum to 0 due to alternating signs stemming from the continuity equation.
@@
In the basis truncated to support $r$, some of the terms needed for cancellation are missing, resulting in an almost-conservation. Let us denote the truncated $M^{2}$ by $Q^{(r)}$, explicitly $Q^{(r)}=\sum_{i,k,|k-i|<r}\sigma_{i}^{z}\sigma_{k}^{z}$. In particular, the $j_{i}j_{k}$ terms coming from $\mathcal{U}(\sigma_{i}^{z}\sigma_{k}^{z})$ with $\left\lvert k-i\right\rvert=r-1$, that would cancel out with the same (but negative) terms from $\mathcal{U}(\sigma^{z}_{i}\sigma^{z}_{k+1})$, do not cancel out, since $\sigma^{z}_{i}\sigma^{z}_{k+1}$ is not included in $Q^{(r)}$ (and similarly for $\sigma_{i}^{z}j_{k}$ terms). Importantly, one can check that for $\left\lvert k-i\right\rvert<r-1$ the cancellation still occurs and all such terms are conserved. The upshot is that the only terms differing between $Q^{(r)}$ and $\mathcal{U}(Q^{(r)})$ are those coming from $\sigma_{i}^{z}\sigma_{k}^{z}$ with $\left\lvert k-i\right\rvert=r-1$, the number of which does not scale with $r$.
The number of terms in $Q^{(r)}$ is equal to the number of indices with |ki|<r|k-i|<r, i.e., it scales as $\sim r$ ($\sim Nr$ in a system of length $N$). On the other hand, the number of term differing between $\mathcal{U}(Q^{(r)})$ and $Q^{(r)}$ is constant, $\sim c$. In the limit $r\to\infty$ the relative difference in small. Therefore, this suggests that $Q^{(r)}$ is almost an eigenvector with eigenvalue being close to $1$, which we estimate to be [^5] More exactly, given a diagonalizable matrix $M$ and any vector $v$, we have $\left\lVert Mv\right\rVert=\left\lVert P^{-1}\Lambda Pv\right\rVert\leq\left\lvert\lambda_{v}\right\rvert\left\lVert P\right\rVert_{\mathrm{op}}\left\lVert P^{-1}\right\rVert_{\mathrm{op}}\left\lVert v\right\rVert$. Here, $\Lambda$ is the diagonal matrix of eigenvalues, $P$ the transition matrix into the eigenbasis of $M$, $\left\lVert\bullet\right\rVert_{\mathrm{op}}$ the operator norm corresponding to $\left\lVert\bullet\right\rVert$ and $\lambda_{v}$ the largest eigenvalue (by magnitude) such that the corresponding eigenvector overlaps with $v$. The inequality gives $\left\lvert\lambda_{v}\right\rvert\geq\frac{1}{\left\lVert P\right\rVert_{\mathrm{op}}\left\lVert P^{-1}\right\rVert_{\mathrm{op}}}\frac{\left\lVert Mv\right\rVert}{\left\lVert v\right\rVert}$, which differs from Eq. (20) only by the factor of $\frac{1}{\kappa(M)}=\frac{1}{\left\lVert P\right\rVert_{\mathrm{op}}\left\lVert P^{-1}\right\rVert_{\mathrm{op}}}$. $\kappa(M)$ is the spectral condition number and measures the non-normality of a matrix (for normal matrices $\kappa=1$, for non-normal $\kappa>1$). For the truncated propagator one can numerically check that $\kappa$ is large and increases with $r$, making the bound uninformative. Nevertheless, the numerical evidence in Fig. 7 suggests that the informal estimate (20) still roughly holds. This can be traced back to the empirical fact that the subleading eigenvector is well behaved and almost equal to $Q^{(r)}$.
@@
$\lvert\lambda\rvert\approx 1-\frac{\left\lVert\mathcal{U}\!\left(Q^{(r)}\right)-Q^{(r)}\right\rVert}{\left\lVert Q^{(r)}\right\rVert}=1-\frac{c}{r}.$ (20)
@@
4.2 Hydrodynamic tails
In systems with conserved quantities, certain correlation functions related to transport exhibit power-law decays, i.e., the so-called hydrodynamic tails [7, 30], due to slow power-law spreading of inhomogeneities of the conserved density (distance and time scaling as $x^{z}\sim t$). A typical example is the correlation function of local magnetization (15), i.e., the Green’s function of the diffusion equation, decaying as $\sim 1/t^{1/2}$ at $x=0$. Similar power-law tails occur also in higher-order corrections to diffusion equation [7, 30], more recently studied in the context of effective field theories [5, 25, 3, 28].
Power-law decay $C(t)\sim t^{-\alpha}$, for some $\alpha>0$, being slower than exponential, cannot be governed by a discrete RP resonance. If we nonetheless assume it expands over a spectrum of eigenvalues as $C(t)=\sum_{j}c_{j}\lambda_{j}^{t}$, where $c_{j}$ are expansions coefficients [cf. Eq. (11)], it must expand over a continuum of eigenvalues (an “RP continuum”) with closing gap to $\lambda=1$ [^6] The argument can be illustrated more explicitly by writing the decomposition as an integral $C(t)=\sum_{j}c_{j}{\rm e}^{-\nu_{j}t}=\int_{0}^{\infty}c(\nu){\rm e}^{-\nu t}\,\mathrm{d}\nu$, where $\lambda={\rm e}^{-\nu}$, $c(\nu)$ is the continuous version of the expansion coefficients and we assumed $\lambda$ to be real for simplicity. In other words, $C(t)$ is the Laplace transform of $c(\nu)$. By employing the inverse Laplace transform, we see that $C(t)=t^{-\alpha}$ is obtained for $c(\nu)=\frac{1}{\Gamma(\alpha)}\nu^{\alpha-1}\theta(\nu)$, where $\Gamma$ is the Gamma function and $\theta$ is the Heaviside step function. We observe that $c(\nu)\neq 0$ for all $\nu>0$ – an RP continuum over all $\lambda\in[0,1]$.. This is precisely what we observed for the correlation function of magnetization in momentum space (16), with the gap closing at $k=0$. Additionally, the quasi-momentum dependence of the RP continuum was exactly determined in that case, $\left\lvert\lambda(k)\right\rvert={\rm e}^{-Dk^{2}}$. The expansion over eigenvalues for all $k$ is expected only for correlation functions of local observables. For extensive observables, the whole continuum must lie in one $k$ sector. A potential RP continuum governing a power-law decay of a $k=0$ extensive observable is depicted in the diagram in Fig. 8.
@@ 4.2 Hydrodynamic tails
Based on the above arguments, we conjecture that there is an RP continuum of eigenvalues for all $|\lambda|$ below the leading RP resonance governing the transport of magnetization. In other words, all $|\lambda|\leq{\rm e}^{-Dk^{2}}$ are part of an RP continuum.
4.2.1 Extensive spin current
@@
where $j_{i}$ is the current operator between sites $i$ and $i+1$, and we introduced the magnetization acting on sites $i$ to $l$ as $M_{[i,l]}\coloneq\sum_{j=i}^{l}\sigma^{z}_{j}$. The closed-form expression for $j_{i}$ is complicated to derive analytically for our 3-layer brickwall circuit, but can be obtained numerically for a given gate by evaluating the LHS of Eq. (21) in a big enough finite circuit. Current densities $j_{i}$ are different for different values of $i$ modulo $3$, and have support on at most $9$ sites. For a 2-layer brickwall circuit (e.g., Fig. 2a) the expressions are more manageable (2 or 4 site operators) and are written out for any magnetization-conserving gate in Ref. [50].
The asymptotics at large $t$ of the local spin current correlation function can be deduced from the local magnetization correlation function and the continuity equation. The diffusive local magnetization correlation function (15) (i.e., the diffusion equation Green’s function) implies
@@
because it appears in the Green-Kubo equation for the diffusion constant [19] [^8] For discrete-time dynamics, the Green-Kubo formula reads $D=\lim_{t_{\mathrm{max}}\to\infty}\lim_{N\to\infty}\frac{1}{N}\big(\frac{1}{2}\left\langle J(0)J(0)\right\rangle+$ $\sum_{t=1}^{t_{\mathrm{max}}}\left\langle J(t)J(0)\right\rangle\big)$. In the considered circuits, it gives $D\approx 0.31$ in the circuit with gate 1 and $D\approx 0.44$ in the circuit with gate 2, which matches our predictions with other methods up to accessible precision (see Fig. 4).. Using Eq. (22) and the fact that the equilibrium correlation function is translationally invariant in space, we get for the leading order of the extensive current in the thermodynamic limit
@@
The long-time prediction coming from the diffusive (i.e., first order) local magnetization correlation function, thus gives exactly 0, which can be traced back to the fact that $\left\langle j(x,t)j(0,0)\right\rangle$ is a total derivative in $x$ [see Eq. (22)] [^9] Note that, interestingly, according to the Green-Kubo formula, the diffusion constant $D$ must acquire its nonzero value at times before the local magnetization correlation function (15) gets its Gaussian form and $\left\langle J(t)J(0)\right\rangle$ becomes zero.. This is very interesting as it means that the leading asymptotic behavior of the correlation function of the extensive current is not $\sim t^{-3/2}$ (in 1d), but is rather given by subleading corrections to the local magnetization correlation function. Correlation function of $J$ is, therefore, a sensitive probe of subleading corrections to diffusion, that would perhaps be hard to see in other correlation functions that are nonzero already in the leading order.
Figure 9: Infinite temperature autocorrelation of extensive spin current. The correlation functions are calculated in circuits with gate 3 with a different number of sites NN. Dashed lines show power-law fits.
Numerical calculation of the extensive current correlation function for one circuit realization is shown in Fig. 9. It decays as a power law for any finite system size $N$, however, the power increases with $N$. It is not clear what will happen in the thermodynamic limit; its power-law dependence may converge to some finite power or the decay may become exponential.
In addition to the inconclusive numerics, it is also not clear what is theoretically expected. Power-law decay $\sim t^{-3/2}$ is expected in systems with an additional conserved quantity, e.g., in Hamiltonian systems [30], and can be explained by the cross transport of the conserved quantities. In systems with only one U(1) conserved quantity, a power-law decay with a higher power was predicted [28], although exponential decay is also sometimes observed, e.g., in the kicked Ising model [47]. The power law predictions are based on the corrections to the diffusive local magnetization correlation function calculated from the effective field theory of diffusion [3], however, from the calculation in Ref. [28], it is not clear what is the first nonzero correction contributing to the extensive current correlation function. Furthermore, subleading corrections can depend on unknown model-dependent parameters.